OpenStudy (anonymous):

Solve the equation: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (anonymous):

-52

OpenStudy (uri):

@Tahamohammed Don't you give direct answers.

OpenStudy (anonymous):

We're not allowed?

OpenStudy (uri):

@Taniphaa Let's do it together..

OpenStudy (anonymous):

@uri We aren't allowed?

OpenStudy (uri):

We have to guide students to the answer.Not just give an answer @Tahamohammed

OpenStudy (anonymous):

Oh I'm sorry I'm new

OpenStudy (uri):

It's alright!

OpenStudy (uri):

First we will solve the brackets, −4(x + 10) − 6 = −3(x − 2) -4x-40-6=-3x+2 Did you get till here?

OpenStudy (jasmineflvs):

what i'm giving the answer!!!

OpenStudy (jasmineflvs):

i did

OpenStudy (jasmineflvs):

happy